# Boosting Ensembles for Statistics of Tails at Conditionally Optimal Advance Split Times

### v1

@@ 5.5 AST selection criteria: aggregate results
Fig. 9c goes beyond the case study to show dispersion indicators averaged across all ancestors. The coefficients of determination for linear and quadratic models (panel c.i) are farther apart on average than they are for the case study, the quadratic model enjoying much higher skill especially during the pivotal 10-20 day range when EI and TE tend to maximize (panels c.(ii,iii)). This validates our choice to use the quadratic model. Overall, the EI, TE, global and local correlations (panels c.(ii-v)) are similar on average to the case study, but smoother.
Note, however, that these averaged dispersion indicators are never used directly in AST selection: the COASTs are chosen separately for each ancestor as the maximizer of its own EI or TE, or at the longest AST such that global or local correlation is above $\rho^{\$}$. This nuance is further illustrated in Fig. 10(a,b), where (EI, TE) are plotted as joint functions of AST and input scale. Whereas the heatmaps are averages over ancestors of EI and TE just like Fig. 9c.(ii,iii), the red circles indicate the fraction of ancestors whose EI or TE is maximized at a particular AST for each particular scale. We call the red circle sizes “COAST frequencies”. For example, at $s=0.24$, the mean EI maximizes at $A=14$ days, and that same AST is the most frequent COAST. However, the second-largest circle indicates that $A=20$ days is a close second-most frequent COAST according to EI. At the same scale, the most frequent COASTs according to TE are $A=18$ and 202020. In general, we gather two patterns from Fig. 10(a,b): the average EI and TE values (i) are well-correlated with their corresponding COAST frequencies, and (ii) both change rapidly at small scales but stabilize above $s\approx 0.24$, at which point the input distributions are close enough to uniform over the $W$-disc. This relative stability is reassuring, but we generally prefer smaller noise which disturbs the model dynamics less. To balance these considerations, we select $s=0.24$ as the nominal scale to examine more closely going forward.
@@ 6 Results: Climatological severity distributions
Having explained the construction of conditional distributions, we now aggregate across ancestors using MoCTail and PoPTail estimators and evaluate the skill of each selection rule by the $\chi^{2}$ divergence of the resulting climatological distribution from ground truth. We first restrict attention to extremes at $y_{0}=\frac{26}{64}L$ and then assess a broader swath of latitudes.
@@
First, consider the simplest AST selection rule $A=A^{\$}$, a uniform AST over all ancestors. We have no a priori principle for $A^{\$}$, so we search through all possible values from 2 to 40 days. Fig. 10c displays the resulting $\chi^{2}$ divergence between the MoCTail and ground truth, as a function of $A^{\$}$ and input scale. A clear optimum emerges at $A^{\$}=14$ days and persists for all scales $s\gtrsim 0.24$, after rapid changes across smaller scales. Red contours also indicate the local correlation, averaged across ancestors to give a smooth and monotonic function of AST. In terms of correlation, the COAST $A^{\$}=14$ days corresponds to $\rho^{\$}\approx 0.92$ depending on the scale, which is slightly above the nominal value $1-(\frac{3}{8})^{2}=0.86$, meaning one should split a little bit closer to the event than the rule of thumb implies.
@@
Overall, the $\chi^{2}$ landscape (inverted) roughly aligns with the EI and TE landscapes, as do their respective optima. This is remarkable and encouraging: allowing each ancestor to determine its own COAST independently in a “liberated” policy, with no knowledge of the ground truth or even other ancestors’ COASTs, leads to a similar solution as the heavy-handed policy of synchronizing them all.
@@
Figure 10: Three optimization landscapes as joint functions of AST and input scale: (a) expected improvement (EI), (b) thresholded entropy (TE), and (c) $\chi^{2}$ divergence between the MoCTail and ground truth. Brighter colors indicate better performance—smaller $\chi^{2}$ divergence or larger EI and TE—and the corresponding “best” ASTs consistently fall in the interior of the domain, across all scales. Contours of local correlation $\rho[c(y_{0},\cdot)]$ are overlaid in (c), giving roughly equivalent correlation levels for any given AST and scale. Red circles in (a,b) indicate the “COAST frequency”: the fraction of ancestors whose (EI, TE) is maximized at the corresponding AST while holding the scale fixed. Note the multiple local maxima in mean AST (brightness), each of which is the global maximum for some significant set of ancestors.
Fig. 11 makes a tail-to-tail comparison between all the AST selection rules, fixing the scale to $s=0.24$ and (in the case of $A^{\$}$ and $A^{\text{\textcent}}$) selecting post-hoc the best-performing threshold ($A^{\$}$ and $\rho^{\$}$ respectively) to set the COASTs. All the rules ($A^{\$},A^{\text{\textcent}},A^{\pounds}$) successfully convert the short DNS tail (left), from which all boosted ensemble members emanate, into a longer tail that tracks closer to the ground truth farther into the extreme severity range. This is borne out visually in the top row, and quantitatively by the consistent reduction in $\chi^{2}$ in the bottom panel across all rules. However, some rules are better than others. Among the “synchronized” COASTs, the constant-AST rule $A^{\$}$ is better than both versions of $A^{\text{\textcent}}$, and the global-correlation version of $A^{\text{\textcent}}$ is a better indicator than the local-correlation version. All three have asymmetric uncertainty bands indicating a large risk of underestimating the ground-truth probabilities. In contrast, both “liberated” COASTs $A^{\pounds}[\text{EI}],A^{\pounds}[\text{TE}]$ produce accurate point estimates and narrow, symmetric uncertainty bands.
Some subtle differences between MoCTail (crosses) and PoPTail (circles) estimates are also visible: whereas PoPTail is more accurate at $A^{\$}$ and both $A^{\text{\textcent}}$s, MoCTail is more accurate when using $A^{\pounds}$. However, one shouldn’t split hairs over small differences in $\chi^{2}$, which might arise from inconsequentially tiny fluctuations at the very upper tail. The more important point is that all rules deliver an improvement over short DNS, and they do so by finding COASTs lying strictly between the shortest and longest options. The actual $A^{\$}$ and $\rho^{\$}$ thresholds are displayed above panels a.(i-vi): “AST =14(8)=14(8)= 14 ( 8 )” means that 14 and 8 are the respective COASTs for MoCTail and PoPTail estimates respectively. By comparing with Fig. 10c, we recognize 14 and 8 as the primary and secondary minima of the $\chi^{2}$ landscape, which also correspond to the local-correlation values 0.98,0.96$\sim$ 0.98,0.96∼ 0.98 , 0.96, which are approximately the optimal $\rho^{\$}$ values noted above Fig. 11a.(iii) (but in reverse order).
@@
Figure 11: CCDF approximations by various mixing criteria and associated errors, at the latitude $y_{0}=\frac{26}{64}L$ and input scale choice $s=0.24$. (a.i) Tail CCDFs according to the long DNS (dashed black line), GPD fit to long DNS (gray line), short DNS (thick black line), and 90% error bar obtained by quantiles over longitudes (gray shading). The ratio of the two CCDFs is shown in a.vii, where deviation from vertical means larger error at a certain severity level, while the $\chi^{2}$ divergence plotted at the left of (c) in black indicates an integrated form of the error. (a.ii) Tail CCDF estimated by boosting ancestors at a fixed AST of 14 (8) days for the MoCTail (PoPTail) estimators, shown in a purple solid line with crosses (dotted purple line with crosses), overlaid on the ground truth. The specific AST values are chosen to best match the ground truth according to $\chi^{2}$ divergence. Because this requires ground truth knowledge, the $\chi^{2}$ divergences must be interpreted as practical lower bounds. The 90% error bar applies to the MoCTail estimator only, and comes from bootstrapping on entire “families” or in other words mixture components (not individual descendants) and choosing the best AST (by the $\chi^{2}$ divergence) for each particular subsample. The error bar widths, too, must then represent lower bounds. Panels a.(iii,iv) show the analogous tail approximations using thresholds of (local, global) correlations as AST selection criteria. Panels a.(v,vi) show the tail approximations obtained by maximizing (EI, TE), which unlike the other criteria do not rely on knowing the ground truth to select ancestor-wise ASTs. All ratios with the ground truth CCDF are overlaid in (b), and all $\chi^{2}$ divergences are shown in (c).
@@
Similar patterns hold across target latitudes, but with some notable caveats. The $\chi^{2}$ divergences of each selection rule are plotted in Fig. 12, of which Fig. 11c is one slice. The most obvious and important point holds: perturbed ensembles improve upon the baseline short-DNS estimate, for almost all latitudes and AST selection rules. The coordinated selection rules ($A^{\$}$ and $A^{\text{\textcent}}$) are the most reliable, and $A^{\pounds}[\text{TE}]$ is slightly less so, but in our opinion is still justified by its “liberated” quality. But $A^{\pounds}[\text{EI}]$ is far less reliable; its favorable performance noted above in Fig. 11 is peculiar to the latitude $y_{0}=\frac{26}{64}L$. At other latitudes, especially in the upper half of the domain, it is similar or worse in skill than short DNS. Even so, it tends to fail by overestimating severities, which we have confirmed by examining the corresponding CCDFs (not shown), and thus it may serve as a useful upper bound.
@@
Figure 12: Performance of all AST selection criteria, measured by $\chi^{2}$ divergence, across all latitudes. Box radius and input scale are the same as in Fig. 11. Black line and gray envelope represent the error from the short DNS and its 90% error bar according to quantiles across longitudes. Panels a-e parallel Fig. 11a.(ii-vi). Solid lines and crosses represent the MoCTail estimator, while dotted lines with open circles represent the PoPTail estimator. Panel (f) displays the bottom topography for reference, which seems to correlate roughly with the KL divergences shown.
@@
The various estimators and AST selection rules have differences in skill, but a more important commonality: all of them indicate that an optimal advance split time exists, somewhere strictly between zero and infinity, which is not a foregone conclusion. Fig. 10 shows clear intermediate optima when targeting the single latitude $y_{0}=\frac{26}{64}L$, and Fig. 13 extends this result to all latitudes by stacking together cross-sections of the per-latitude counterparts of Fig. 10 at $s=0.24$. The COAST frequency and mean-TE landscapes have broad ridges that meander slowly in AST space with latitude, approximately in phase with topography: smaller ASTs are favored at $y_{0}\approx\frac{26}{64}L$, where topography is minimized and meridional wind shear is negative, and larger ASTs are favored at $y_{0}\approx\frac{38}{64}L$, where topography is maximized and meridional wind shear is positive. A similar pattern, but with bigger swings, is seen in the $\chi^{2}$ landscape. All these patterns are a bit noisy, especially for the COAST frequencies and $\chi^{2}$-COAST locations, since both come from an inherently unstable “argmax” function. Nonetheless, the detailed latitude dependence is only a secondary effect on top of the main point, which is clearly demonstrated: splitting is most effective at intermediate ASTs rather than very short or long ASTs.
@@
We can also now evaluate the $\frac{3}{8}$ rule from Finkel and O’Gorman, (2024) in this broader multi-latitude context, though here we simplify the procedure by first averaging $\rho$ across ancestors and then calculating $A^{\$}$ as a threshold-crossing time of that average, which we call $A_{3/8}^{\$}$, rather than averaging times $A_{n}^{\text{\textcent}}[\rho^{\$}=1-(\frac{3}{8})^{2}]$ across ancestors. The same conclusion holds either way. The AST values $A_{3/8}^{\$}$ are overlaid on the $\chi^{2}$ heatmap (Fig. 13d) as blue curves. The solid curve, representing a level set of ancestor-averaged global correlation, should be constant with latitude and varies only due to sampling errors. Likewise, the dashed curve, representing a level set of ancestor-averaged local correlation, should be symmetric with respect to latitude because of the symmetries in tracer dynamics, as should all the level sets in panel c. Since the $A^{\$}$ varies differently with latitude, exhibiting roughly odd symmetry about the midline, the $\frac{3}{8}$ rule cannot possibly be optimal for all latitudes simultaneously. More fundamentally, the COAST depends on more than just a generic metric for ensemble dispersion: it must also depend on the features of the tail being sampled, which in this case is the only possible source of broken symmetry (see Fig. 4).
@@
However, both versions of $A_{3/8}^{\$}$ run right through the mean position of the meandering $\chi^{2}$ valley and associated COASTs, performing about as well as any such highly-constrained synchronized $A^{\$}$ could do. Thus, the $\frac{3}{8}$ rule retains its relevance as a starting point for more refined optimization more tailored to the event, at least for this QG system. Whether the $\frac{3}{8}$ rule generalizes further to more heterogeneous systems as the “optimal synchronized AST” remains to be seen.
@@
Figure 13: Optimization landscapes and optimal ASTs across latitudes, again fixing the box half-width to $\frac{2}{64}L$ and the input scale to $s=0.24$. (a) Frequencies of conditionally optimal ASTs (COASTs), in the maximum-thresholded entropy sense, at each latitude. E.g., at $y_{0}/L=26$, the two adjacent bright pixels at AST =18,20=18,20= 18 , 20 indicate that for a large fraction of ancestors, the highest-entropy descendant ensemble is the one launched 18 or 20 days in advance of the peak (pixel brightness has the same meaning here as circle size in Fig. 10). (b) Thresholded entropy as a function of AST, also normalized to (0, 1) at each latitude, with absolute ranges shown in (c). This landscape is smoother than $\chi^{2}$ and varies less dramatically with latitude, but exhibits directionally similar trends. (d) $\chi^{2}$ divergence as a function of AST and latitude, normalized to the range 0-1 (white-black) separately at each latitude (see the lower and upper bounds in (e)) so that different latitudes are visually comparable. Red crosses mark the optimal AST at each latitude. Cyan (solid, dashed) curves mark the AST at which the (global, local) correlations, averaged across ancestors, reach $1-(\frac{3}{8})^{2}$. This nominal choice is based on Finkel and O’Gorman, (2024), and falls squarely in the middle of the latitude-dependent ASTs. (f) Contour map of local correlation, averaged over ancestors, as a function of AST and latitude. The levels range from 0.22 (left-most dotted black curve, fragmented by boundary) to 0.99 (rightmost solid black curve), evenly spaced in a stretched sigmoid scale (levels are not shown and are shown only for qualitative purposes). The reference level $1-(\frac{3}{8})^{2}$ appears dashed in cyan. (g) Bottom topography for reference.

### v2

@@ 5.5 AST selection criteria: aggregate results
Fig. 11 goes beyond the case study to show dispersion indicators averaged across all ancestors. The coefficients of determination for linear and quadratic models (Fig. 11a) are farther apart on average than they are for the case study (see Fig. 8e), the quadratic model enjoying much higher skill especially during the pivotal 10-20 day range when EI and TE tend to maximize (Fig. 11b,c). This validates our choice to use the quadratic model. Overall, the EI, TE, global and local correlations (Fig. 11 b-e) are similar on average to the case study, but smoother.
Figure 11: Ensemble dispersion metrics averaged across ancestors at $y_{0}=26/64L$. (a) Coefficients of determination for linear (cyan) and quadratic (orange) regressions, averaged across ancestors. (b-e) same quantities as in in Fig. 10(a-d) but averaged across ancestors, with only the largest and smallest scales shown (red: $s=0.06$, blue: $s=0.9$). Shaded regions show the areas between truncated upper and lower means. E.g., for the correlation $\rho$, the truncated upper mean is the mean of $\rho$ across ancestors with above-average $\rho$: $\mathbb{E}[\rho|\rho>\mathbb{E}[\rho]]$, separately at each AST. We choose truncated means as a compromise between quantiles (which are erratic for the relatively small sample size of ancestors) and standard deviation envelopes (which can misleadingly fall outside the bounds $[0,1]$ to which $\rho$ is constrained).
Note, however, that these averaged dispersion indicators are never used directly in AST selection: the COASTs are chosen separately for each ancestor as the maximizer of its own EI or TE, or at the longest AST such that global or local correlation is above $\rho^{\text{U}}$. This nuance is further illustrated in Fig. 12(a,b), where (EI, TE) are plotted as joint functions of AST and input scale. Whereas the heatmaps are averages over ancestors of EI and TE just like Fig. 9c.(ii,iii), the red circles indicate the fraction of ancestors whose EI or TE is maximized at a particular AST for each particular scale. We call the red circle sizes “COAST frequencies”. For example, at $s=0.24$, the mean EI maximizes at $A=14$ days, and that same AST is the most frequent COAST. However, the second-largest circle indicates that $A=20$ days is a close second-most frequent COAST according to EI. At the same scale, the most frequent COASTs according to TE are $A=18$ and $20$. In general, we gather two patterns from Fig. 12(a,b): the average EI and TE values (i) are well-correlated with their corresponding COAST frequencies, and (ii) both change rapidly at small scales but stabilize above $s\approx 0.24$, at which point the input distributions are close enough to uniform over the $W$-disc. This relative stability is reassuring, but we generally prefer smaller noise which disturbs the model dynamics less. To balance these considerations, we select $s=0.24$ as the nominal scale to examine more closely going forward.
@@ 6 Results: Climatological severity distributions
Having explained the construction of conditional distributions, we now aggregate across ancestors using MoCTail and PoPTail estimators to obtain our estimates of the climatological severity distribution from the boosted ensembles. We evaluate the skill of each AST selection rule by the $\chi^{2}$ divergence of the resulting climatological distribution from ground truth as obtained from the long DNS. We first restrict attention to extremes at $y_{0}=\frac{26}{64}L$ and then assess a broader swath of latitudes.
@@
First, consider the simplest AST selection rule $A=A^{\text{U}}$, a uniform AST over all ancestors. We have no a priori principle for $A^{\text{U}}$, so we search through all possible values from 2 to 40 days. Fig. 12c displays the resulting $\chi^{2}$ divergence between the MoCTail and ground truth, as a function of $A^{\text{U}}$ and input scale. A clear optimum emerges at $A^{\text{U}}=14$ days and persists for all scales $s\gtrsim 0.24$, after rapid changes across smaller scales. Red contours also indicate the local correlation, averaged across ancestors to give a smooth and monotonic function of AST. In terms of correlation, the COAST $A^{\text{U}}=14$ days corresponds to $\rho^{\text{U}}\approx 0.92$ depending on the scale, which is slightly above the nominal value $1-(\frac{3}{8})^{2}=0.86$, meaning one should split a little bit closer to the event than the rule of thumb implies.
@@
Overall, the $\chi^{2}$ landscape (inverted) roughly aligns with the EI and TE landscapes, as do their respective optima. This is remarkable and encouraging: allowing each ancestor to determine its own COAST independently, with no knowledge of the ground truth or even other ancestors’ COASTs, leads to a similar solution as the policy of synchronizing them all. Boosting based on EI and TE, therefore, is more parallelizable (optimizations are decoupled across ancestors), extensible (new ancestors can be added without changing the optimal split times for pre-existing ancestors), and interpretable (one can see the optimum clearly based on a case study, without complicated averaging procedures across initial conditions).
@@
Figure 12: Three optimization landscapes as joint functions of AST and input scale for $y_{0}=(26/64)L$: (a) expected improvement (EI), (b) thresholded entropy (TE), and (c) $\chi^{2}$ divergence between the MoCTail and ground truth. Lighter gray indicates better performance—smaller $\chi^{2}$ divergence or larger EI and TE—and the corresponding “best” ASTs consistently fall in the interior of the domain, across all scales. Contours of local correlation $\rho[c(y_{0},\cdot)]$ are overlaid in (c), giving a rough map of correspondence between correlation levels and AST. The size of red circles in (a,b) indicate the “COAST frequency”: the fraction of ancestors whose (EI, TE) is maximized at the corresponding AST while holding the scale fixed. Note the multiple local maxima in mean EI and TE (as indicated by the lightness of the gray color in (a,b)), each of which is the global maximum for some significant set of ancestors.
Fig. 13 makes a tail-to-tail comparison between all the AST selection rules (a.i-v: $A^{\text{U}},A^{\text{PC}}$ local and global, $A^{\text{EI}},A^{\text{TE}}$), fixing the scale to $s=0.24$ and (in the case of $A^{\text{U}}$ and $A^{\text{PC}}$) selecting post-hoc the best-performing threshold to set the COASTs. We used subsets of only 11 of the 32 ancestors, resampling such subsets 64 times to obtain medians (solid) and interquartile ranges (shading) on CCDFs. The numerical values of optimal AST and $\rho$ reported above a.(i-iii), with PoPTail optima parenthesized, are the optima obtained from $N=32$, i.e., the best estimates of the true optima; they don’t necessarily correspond to the values used for plotting with $N=11$, which are optimized separately for each resampling. The brown CCDF in panel (a.vi) is the estimate from the unboosted acestors alone (“equal-NN”), and the black is the estimate from a larger number of ancestors to equal the cost of boosting. The curves underneath in panel (b) show the rate of improvement of $\chi^{2}$ with $N$.
In terms of quantitative improvements in $\chi^{2}$, all the rules considered ($A^{\text{U}},A^{\text{PC}},A^{\text{EI}},A^{\text{TE}}$) improve substantially upon an equal-$N$ DNS and modestly upon an equal-cost DNS. The size of the advantage varies with $N$ in the way that we expect from boosting: substantial improvements with moderate $N$, when the DNS has sampled the attractor broadly but sparsely and extremes are within reach by perturbation. The advantage might diminish if $N$ increases enough for DNS to see those extremes without perturbation, but we haven’t reached that regime yet. MoCTail and PoPTail performances are similar, but not identical: PoPTail seems more suited for threshold-based rules ($A^{\text{U}},A^{\text{PC}}$ local and global in b.(i-iii)), whereas MoCTail seems more suited for optimization-based rules ($A^{\text{EI}},A^{\text{TE}}$ in b.(iv,v)).
We selected $N=11$ to display the full CCDFs in Fig. 13(a) as the middle range of values tried, and where enough equal-size ancestor subsets are available for uncertainty quantification by bootstrapping. When comparing with DNS CCDFs, all five rules successfully extend the short, equal-$N$ DNS tail into a longer tail that tracks closer to the ground truth farther into the extreme severity range. They also all find a larger maximum than even the equal-cost DNS found. However, the threshold-based rules exhibit apparent bias, systematically underestimating probabilities for $R^{*}\gtrsim 0.64$ with asymmetric variabilities, whereas the optimization-based rules are both more accurate and more confident.
The COASTs identified by all rules lie strictly between the shortest and longest ASTs considered. For example, $A^{\text{U}}=14$ according to the MoCTail estimator (using all $N=32$ ancestors). By comparing with Fig. 12c, we recognize 14 as the minimum of the $\chi^{2}$ landscape for $s=0.24$ (and larger scales), with an approximate local-correlation equivalent of 0.98.
@@
Figure 13: CCDF approximations by various mixing criteria and associated errors, at the latitude $y_{0}=\frac{26}{64}L$ and input scale choice $s=0.24$. (a.i-v) Tail CCDFs by various estimates using only $N=11$ ancestors, with lines showing medians and bands showing interquartile ranges across many size-11 subsamples of the total set of 32 ancestors. Lines are medians, and bands are interquartile ranges. Dotted lines with open circles are PoPTails, while solid lines with crosses are MoCTails. Dashed black lines show the ground truth estimate. Panel a(i) shows the tail approximation using a single uniform AST indicated at the top: 14 days for MoCTail and 8 days (parenthesized) for PoPTail. Panels a.(ii,iii) show the tail approximations using thresholds of (local, global) correlations as AST selection criteria. Panels a.(iv,v) show the tail approximations obtained by maximizing (EI, TE), which unlike the other criteria do not rely on knowing the ground truth to select ancestor-wise ASTs, either directly or through threshold choice. (a.vi) also shows estimates from DNS with equal cost to boosting on 11 ancestors (black stars, gray envelope) and DNS from only $N=11$ peaks (brown circles and envelope), in both cases estimating uncertainty by longitudinal rotation. The GPD fit to ground truth is shown as a gray curve. In a.(i-iii), the thresholds shown at the top (PoPTail thresholds parenthesized) are obtained by using all 32 ancestors, but the CCDFs displayed each choose an AST to minimize $\chi^{2}$ divergence from ground truth, separately for each subsample. Because this requires ground truth knowledge, the $\chi^{2}$ divergences must be interpreted as practical lower bounds. The 90% error bar applies to the MoCTail estimator only, and comes from bootstrapping on entire “families” or in other words mixture components (not individual descendants) and choosing the best AST (by $\chi^{2}$ divergence) for each particular subsample. The error bar widths, too, must then represent lower bounds. (b) $\chi^{2}$ values for the estimator directly above in each case as a function of $N$, and compared with DNS at equal cost and equal NN. DNS does not run long enough to equal the total cost accrued by boosting 32 ancestors, so the black curve stops before the others.
@@
Similar patterns hold across target latitudes, but with some notable caveats. The $\chi^{2}$ divergences of each selection rule are plotted in Fig. 14, of which Fig. 13c is one slice. The most obvious and important point holds: perturbed ensembles improve upon the DNS equal-$N$ estimate, for almost all latitudes and AST selection rules, and they also improve on the equal cost estimate in many cases. But $A^{\text{EI}}$ is less reliable; its favorable performance noted above in Fig. 13 is peculiar to the latitude $y_{0}=\frac{26}{64}L$. At some other latitudes, it is similar or worse in skill than equal-$N$ and even equal-cost DNS. Even so, it tends to fail by overestimating severities, which we have confirmed by examining the corresponding CCDFs (not shown), and thus it may serve as a useful upper bound. The MoCTail and PoPTail estimators are similar in quality across latitudes, but as observed in Fig. 13, PoPTail has an advantage with threshold-based rules ($A^{\text{U}},A^{\text{PC}}$ local and global) whereas MoCTail performs better with optimization-based rules ($A^{\text{EI}},A^{\text{TE}}$).
@@
Figure 14: Performance of all AST selection criteria, measured by $\chi^{2}$ divergence, across all latitudes for $s=0.24$ and $N=10$ or $1$, whichever is nearest to 1/3 the number of ancestors found for the latitude in question (sometimes less than 32). Black line and gray envelope represent the error from the short DNS and its 90% error bar according to quantiles across longitudes. Panels a-e parallel Fig. 13a.(ii-vi). Solid lines and crosses represent the MoCTail estimator, while dotted lines with open circles represent the PoPTail estimator.
@@
The various estimators and AST selection rules have differences in skill, but a more important commonality: all of them indicate that an optimal advance split time exists that is strictly positive, which is not a foregone conclusion in light of standard rare event algorithms like adaptive multilevel splitting (AMS; Lestang et al., 2018) without “trying early”. Fig. 12 shows clear intermediate optima when targeting the single latitude $y_{0}=\frac{26}{64}L$, and Fig. 15 extends this result to all latitudes by stacking together cross-sections of the per-latitude counterparts of Fig. 12 at $s=0.24$. The COAST frequency and mean-TE landscapes have broad ridges that meander slowly in AST space with latitude, approximately in phase with topography: smaller ASTs are favored at $y_{0}\approx\frac{26}{64}L$, where topography is minimized and meridional wind shear is negative, and larger ASTs are favored at $y_{0}\approx\frac{38}{64}L$, where topography is maximized and meridional wind shear is positive. A similar pattern, but with bigger swings, is seen in the $\chi^{2}$ landscape. All these patterns are a bit noisy, especially for the COAST frequencies and $\chi^{2}$-COAST locations, since both come from an inherently unstable “argmax” function. Nonetheless, the detailed latitude dependence is only a secondary effect on top of the main point, which is clearly demonstrated: splitting is most effective at intermediate ASTs rather than very short or long ASTs.
@@
We can also now evaluate the $\frac{3}{8}$ rule from Finkel and O’Gorman (2024) in this broader multi-latitude context, though here we simplify the procedure by first averaging $\rho$ across ancestors and then calculating $A^{\text{U}}$ as a threshold-crossing time of that average, which we call $A_{3/8}^{\text{U}}$, rather than averaging times $A_{n}^{\text{PC}}[\rho^{\text{U}}=1-(\frac{3}{8})^{2}]$ across ancestors. The same conclusion holds either way. The AST values $A_{3/8}^{\text{U}}$ are overlaid on the $\chi^{2}$ heatmap (Fig. 15d) as blue curves. The solid curve, representing a level set of ancestor-averaged global correlation, should be constant with latitude and varies only due to sampling errors. Likewise, the dashed curve, representing a level set of ancestor-averaged local correlation, should be approximately symmetric with respect to latitude because of the symmetric tracer boundary conditions and approximate mirror symmetry in velocities, as should all the level sets in panel c. Since the $A^{\text{U}}$ varies differently with latitude, exhibiting roughly odd symmetry about the midline, the $\frac{3}{8}$ rule cannot possibly be optimal for all latitudes simultaneously. More fundamentally, the COAST depends on more than just a generic metric for ensemble dispersion: it must also depend on the features of the tail being sampled, which in this case is the only possible source of broken symmetry (see Fig. 4).
@@
However, both versions of $A_{3/8}^{\text{U}}$ run right through the mean position of the meandering $\chi^{2}$ valley and associated COASTs, performing about as well as any such highly-constrained synchronized $A^{\text{U}}$ could do. Thus, the $\frac{3}{8}$ rule retains its relevance as a starting point for more refined optimization more tailored to the event, at least for this QG system. Whether the $\frac{3}{8}$ rule generalizes to more heterogeneous systems as the “optimal synchronized AST” requires further investigation. We found it provides some guidance for temperature and precipitation extremes in an idealized general circulation model, but overestimated the optimal AST in both cases (Finkel and O’Gorman, 2025).
@@
Figure 15: Optimization landscapes and optimal ASTs across latitudes, again fixing the input scale to $s=0.24$. (a) Frequencies of conditionally optimal ASTs (COASTs), in the maximum-thresholded entropy sense, at each latitude, with whiter shading indicating higher frequency. E.g., at $y_{0}/L=26/64$, the two adjacent bright pixels at AST =18,20=18,20 indicate that for a large fraction of ancestors, the highest-entropy descendant ensemble is the one launched 18 or 20 days in advance of the peak. (b) Thresholded entropy as a function of AST, normalized to the range 0-1 (black-white, so brighter is better) separately at each latitude. This landscape is smoother than $\chi^{2}$ and varies less dramatically with latitude, but exhibits directionally similar trends. (c) $\chi^{2}$ divergence as a function of AST and latitude, normalized to the range 0-1 (white-black, so brighter is better) separately at each latitude so that different latitudes are visually comparable. Red crosses mark the optimal AST at each latitude. Cyan (solid, dashed) curves mark the AST at which the (global, local) correlations, averaged across ancestors, reach $1-(\frac{3}{8})^{2}$. This nominal choice is based on Finkel and O’Gorman (2024), and falls squarely in the middle of the latitude-dependent ASTs. (d) Contour map of local correlation, averaged over ancestors, as a function of AST and latitude. The levels range from 0.22 (left-most dotted black curve, fragmented by boundary) to 0.99 (rightmost solid black curve), evenly spaced in a stretched sigmoid scale (levels are shown and only for qualitative purposes). The reference level $1-(\frac{3}{8})^{2}$ appears dashed in cyan. (g) Bottom topography for reference.